OpenStudy (deerhunter15):

Use the Babylonian method to approximate sqrt 19  to the nearest hundredth.

OpenStudy (misty1212):

hi!!

OpenStudy (deerhunter15):

Hello

OpenStudy (misty1212):

first make a guess. i guess 4, which is not too close, because \(4^2=16\) but close enough

OpenStudy (misty1212):

then for the "babylonian"method the next guess is the average of 4 and 19/4 i.e. \[\frac{1}{2}(4+\frac{19}{4})\]

OpenStudy (misty1212):

i get \[\frac{35}{8}\] or \[4.375\]

OpenStudy (misty1212):

for the next one, you repeat the process, only with \(4.375\) instead of 4 i.e. \[\frac{1}{2}(4.375+\frac{19}{4.375})\] at this point you want a calculator

OpenStudy (misty1212):

then do it again etc should be correct to two decimal places in only 3 tries or so